# Projection effects in barred galaxies causing wrong interpretation of radial flows

### v1

@@ 6.2 Comparison with previous studies
In a large HI survey, Di Teodoro & Peek (2021) found that the strongest apparent radial velocities were always associated with bars or spiral arms, i.e. non-axisymmetric structures. Yet, the measured velocities were low ($|V_{\mathrm{rad}}|\lesssim 10$ km s-1), typically three to seven times lower than those obtained here. Our results confirm that bars systematically increase the uncertainty in $V_{\mathrm{rad}}$ and can easily mimic radial flows.
In this work, we used a simulated galaxy but several observed galaxies that have been modeled with 3D-Barolo show the same pattern: apparent inflows or outflows in barred regions are often projection artifacts.
NGC 1808, a strongly barred late-type spiral galaxy (de Vaucouleurs et al., 1991), is known for having large-scale outflows perpendicular to the disk, due to star formation (Salak et al., 2016). Audibert et al. (2021) derived CO kinematics with 3D-Barolo and found large radial velocities in the central 500 pc, where a nuclear spiral is present. From 3D-Barolo’s interpretation, there was an apparent inflow. An inflow was confirmed by gravitational torque analysis, which showed negative torques consistent with true inflow, but with lower amplitude. At smaller scales, they identify the variations of the PA as perturbations caused by the decoupling of the torus. In our case, the scale is too large to be affected by nuclear structures like this, but this result shows that a changing PA could be attributed to many physical phenomena as a warped disk, a bar that has not been identified or a torus at smaller scales. A similar case is the barred galaxy NGC 613, in which Audibert et al. (2019) observed and characterized both a radial inflow through gravity torques, and an outflow perpendicular to the disk.
@@
Dall’Agnol de Oliveira et al. (2025) additionally reported non-circular motions in position–velocity (PV) diagrams along the minor axis (see their Fig. 6), which they attributed to disturbed CO kinematics. In our simulated barred galaxy, similar features arise naturally from the projection of bar-driven elliptical orbits (Fig. 10), without any actual radial inflow or outflow. This suggests that some of the observed signatures may also be consistent with non-axisymmetric streaming motions. Consequently, part of the radial flows reported in the three Seyfert galaxies, could arise from this degeneracy between true radial motions and the LoS projection of elliptical bar orbits.
Figure 10: Position-velocity diagrams along the major axis at $\phi=270^{\circ}$ (top panel) and minor axis at $\phi=0^{\circ}$ (bottom panel) for gSb inclined at $i_{\rm true}=40^{\circ}$ in configuration (a) of Fig. 3. The red contours are the best fit from 3D-Barolo.
@@ 7 Conclusions
Acknowledgment
This work was supported by the Programme National Cosmology et Galaxies (PNCG) of CNRS/INSU with INP and IN2P3, co- funded by CEA and CNES. The data were processed using the Gildas package. We made use of the NASA/IPAC Extragalactic Database (NED), and of the HyperLeda database (http://leda.univ-lyon1.fr). This paper made use of GalMer data: (http://galmer.obspm.fr). This work made use of Astropy (http://www.astropy.org) a community-developed core Python package and an ecosystem of tools and resources for astronomy. We used the algorithm 3D-Barolo (Di Teodoro & Fraternali, 2015; Di Teodoro & Peek, 2021).

### v2

@@ 6.2 Comparison with previous studies
In a large HI survey, Di Teodoro & Peek (2021) found that the strongest apparent radial velocities were always associated with bars or spiral arms, i.e. non-axisymmetric structures. Our results confirm that bars systematically increase the uncertainty in $V_{\mathrm{rad}}$ and can easily mimic radial flows.
Even when the geometrical parameters were set to their true values, the rotation curve estimations were incorrect due to the bar. Similar results were obtained by Liu et al. (2025) using a tilted-ring method without a radial velocity while fixing the inclination at $i=45^{\circ}$ and the PA at $\phi=0^{\circ}$. They found that the rotation curves were strictly dependent on the bar’s orientation in the sky plane. Their simulations at $\Delta\phi=\pm 45^{\circ}$ (see their Fig. 1) respectively match our configurations (c) and (a) of Fig. 3 because the gas rotates clockwise, with $\Delta\phi$ the difference between the galaxy’s PA and the bar’s. In these cases, they identified a ’dip’ feature in the rotation curves within the bar region: a central value higher than the circular velocity followed by a local minimum still within the bar region. We observed the same feature on the rotation curves derived for $30^{\circ}\leq i_{\rm true}\leq 50^{\circ}$ (Fig. 7) in all configurations, unlike Liu et al. (2025) who found a peak when $\Delta\phi$ is negative that we do not retrieve. They also studied the rotation curves of 45 galaxies from the PHANGS-ALMA survey, including 29 barred galaxies. The dip feature was observed in 9 of the 11 galaxies with $|\Delta\phi\leq 40^{\circ}|$. According to Liu et al. (2025), the strength of the bar is related to the strength of the dip. Although the feature is not dominant in the sample of barred galaxies, its absence in the sample of non-barred galaxies shows that it is indeed bar-induced. They suppose this is due to bar orientation. Additionally, identifying such features is complicated by other sources of changes in rotation curves, such as spiral arms or massive bulges within the bar region. Besides that, they proposed a misaligned ellipses model showing how projection creates velocity artifacts that can be misinterpreted as radial flows, as we demonstrated. Because this geometric model alone cannot fully recover the intrinsic circular velocity, they further introduced an empirical, first-order correction to the rotation curve. This correction is based on calibrating the systematic deviations induced by the bar as a function of bar-disk position angle offset and radius.
In this work, we used a simulated galaxy but several observed galaxies that have been modeled with 3D-Barolo show the pattern we highlight in this study: apparent inflows or outflows in barred regions are often projection artifacts.
NGC 1808, a strongly barred late-type spiral galaxy (de Vaucouleurs et al., 1991), is known for having large-scale outflows perpendicular to the disk, due to star formation (Salak et al., 2016). Audibert et al. (2021) derived CO kinematics with 3D-Barolo and found large radial velocities in the central 500 pc, where a nuclear spiral is present. From 3D-Barolo’s interpretation, there was an apparent inflow. An inflow was confirmed by gravitational torque analysis, which showed negative azimuthally-averaged and gas-weighted torques consistent with true inflow, but with lower amplitude. At smaller scales, they identify the variations of the PA as perturbations caused by the decoupling of the torus. In our case, the scale is too large to be affected by nuclear structures like this, but this result shows that a changing PA could be attributed to many physical phenomena as a warped disk, a bar that has not been identified or a torus at smaller scales. A similar case is the barred galaxy NGC 613, in which Audibert et al. (2019) observed and characterized both a radial inflow through gravity torques, and an outflow perpendicular to the disk.
@@
Dall’Agnol de Oliveira et al. (2026) additionally reported non-circular motions in position-velocity (PV) diagrams along the minor axis (see their Fig. 6), which they attributed to disturbed CO kinematics. In our simulated barred galaxy, similar features arise naturally from the projection of bar-driven elliptical orbits (Fig. 10), without any actual radial inflow or outflow. This suggests that some of the observed signatures may also be consistent with non-axisymmetric streaming motions. Consequently, part of the radial flows reported in the three Seyfert galaxies, could arise from this degeneracy between true radial motions and the LoS projection of elliptical bar orbits.
Figure 10: Position-velocity diagrams along the major axis at $\phi=270^{\circ}$ (top panel) and minor axis at $\phi=0^{\circ}$ (bottom panel) for gSb inclined at $i_{\rm true}=40^{\circ}$ in configuration (a) of Fig. 3. The red contours are the best fit from 3D-Barolo obtained with a free PA and a free inclination angle (see Sect. 5.1).
We showed that projection effects can induce a misinterpretation of radial motions in barred spiral galaxies. Identifying non-axisymmetric structures is therefore a prerequisite before interpreting apparent radial motions as genuine inflows or outflows. In particular, true radial inflows of gas are expected to be weak on average. They can be reliably measured through gravitational torque computations or by comparing them with numerical simulations that match observations. The latter approach can in principle provide detailed constraints, but in practice it requires constructing dynamical models that closely reproduce the observed morphology and kinematics, which is challenging. In addition, kinematic methods based on gas flows along bar dust lanes, calibrated using numerical simulations, offer another way to estimate inflow rates (e.g. Sormani & Barnes (2019); Sormani et al. (2023)). These methods assume that, in the frame co-rotating with the bar, the gas velocity vectors are approximately aligned with the dust lanes. They further rely on simulations to calibrate the fraction of gas that overshoots the nuclear ring and join the other bar lane. As a result, the dust lanes approach provides valuable constraints on non-circular motions and on the gas angular-momentum loss. However, translating the kinematics measured along dust lanes into net inflow rates still depends on the extent to which the adopted simulations capture the true gas dynamics of individual galaxies.
@@ 7 Conclusions
Acknowledgements.
This work was supported by the Programme National Cosmology et Galaxies (PNCG) of CNRS/INSU with INP and IN2P3, co- funded by CEA and CNES. The data were processed using the Gildas package. We made use of the NASA/IPAC Extragalactic Database (NED), and of the HyperLeda database (http://leda.univ-lyon1.fr). This paper made use of GalMer data (http://galmer.obspm.fr). This work made use of Astropy (http://www.astropy.org) a community-developed core Python package and an ecosystem of tools and resources for astronomy. We used the algorithm 3D-Barolo (Di Teodoro & Fraternali, 2015; Di Teodoro & Peek, 2021).
Appendix A Plot of the fitted parameters for a free PA and a fixed inclination angle